# Measuring multipartite quantum correlations by thermodynamic work extraction

### v1

@@ Abstract
Quantum correlations are at the core of quantum mechanics and play a crucial role in various fields such as quantum information, condensed matter physics, high energy physics, and quantum thermodynamics. While bipartite quantum correlations have been extensively studied, multipartite quantum correlations in many-body systems remain elusive due to their complex structure. In particular, a primary challenge lies in the fact that the calculation of multipartite quantum correlation measure often requires exponential cost with respect to the system size. In this work, we tackle this problem by adopting a thermodynamic approach; we introduce a measure of multipartite quantum correlations based on the difference in extractable thermodynamic work by global operations and local operations and classical communication (LOCC). This can be regarded as a multipartite generalization of the work deficit, which has attracted attention as a thermodynamic measure of bipartite quantum correlation. A distinguishing feature of the thermodynamic approach to multipartite quantum correlation is that we can compare the degree of quantum correlations with clear operational meaning; for example, it is shown that W state has more quantum benefit in work extraction than GHZ state. Importantly, we develop an efficient calculation method to calculate the multipartite work deficit. This efficient method works for a special class of quantum many-body systems described by matrix product states (MPS), where the numerical cost is shown to be proportional to the system size, significantly reducing the exponential cost required for the direct calculations. We demonstrate this efficient method in representative MPS such as the AKLT state and the cluster state, and analytically obtain the exact values of this measure. We further show that a quantum phase transition described by MPS is well captured by the multipartite work deficit, where a singular behavior appears and becomes sharper under coarse-graining, at the transition point. This shows that the multipartite work deficit does not only highlight the fundamental connection between multipartite quantum correlations and quantum thermodynamics, but also serves as an efficiently-computable probe of the structures of quantum many-body systems including their quantum phase transitions.
@@ I Introduction
In light of recent developments in quantum technologies Gross and Bloch (2017); Schäfer et al. (2020); Kjaergaard et al. (2020); Krantz et al. (2019); Lukin et al. (2020); Altman et al. (2021), quantum correlations are becoming increasingly important not only in quantum computation and quantum communication Gisin and Thew (2007); Nielsen and Chuang (2010); Briegel et al. (2009) but also in various fields of physics including condensed matter physics Amico et al. (2008); De Chiara and Sanpera (2018) and high energy physics Ryu and Takayanagi (2006a, b); Takayanagi (2012). For example, quantum correlations have been utilized as useful probes for understanding critical phenomena Vidal et al. (2003); Calabrese and Cardy (2004, 2009), topological order Kitaev and Preskill (2006); Levin and Wen (2006), and information scrambling Swingle et al. (2016). While the understanding on bipartite quantum correlations has been well-established through various measures with clear operational meanings Plenio and Virmani (2005); Horodecki et al. (2009); Modi et al. (2012); Bera et al. (2017), multipartite quantum correlations remain a very active research avenue Plenio and Vedral (2001); Wei et al. (2004); Barnum and Linden (2001); Wei and Goldbart (2003); Shimony (1995); Orús (2008); Shi et al. (2010); Orús and Wei (2010); Huang and Lin (2010); Mousolou et al. (2013); Orús et al. (2014); Shi et al. (2016); Orus and Wei (2010); Rulli and Sarandy (2011); Campbell et al. (2011, 2013); Power et al. (2015); Sun et al. (2015); Huang et al. (2018); Braga et al. (2012); Xu (2013); Liu et al. (2014); Coffman et al. (2000); Radhakrishnan et al. (2020); Hu et al. (2018); Hyllus et al. (2012); Tóth (2012); Fröwis and Dür (2012); Hauke et al. (2016); Mermin (1990); Ardehali (1992); Pezze et al. (2017), mainly because of their rich and complicated structure Gühne and Tóth (2009); Bengtsson and Zyczkowski (2016). For example, even in tripartite quantum states, there are different types of quantum correlations represented by the W state and the GHZ state.
@@
Furthermore, we develop an efficient calculation method of the multipartite quantum correlations for a special class of quantum many-body systems described by matrix product states (MPS) Perez-Garcia et al. (2006); Verstraete et al. (2008); Cirac et al. (2017, 2021); Orús (2014). While the cost of naive direct calculation of the multipartite work deficit grows exponentially with the system size as is the case for many other measures of multipartite quantum correlations Plenio and Vedral (2001); Wei et al. (2004); Barnum and Linden (2001); Wei and Goldbart (2003); Shimony (1995); Orús (2008); Shi et al. (2010); Orús and Wei (2010); Mousolou et al. (2013); Huang and Lin (2010); Orús et al. (2014); Shi et al. (2016); Orus and Wei (2010); Rulli and Sarandy (2011); Campbell et al. (2011, 2013); Power et al. (2015); Sun et al. (2015); Huang et al. (2018), the cost of our efficient method is proportional to the system size enabling the significant cost reduction (Table. 1). It is noteworthy that this efficient method is not restricted to the work deficit but also is applicable to another measure of multipartite quantum correlation (i.e., global quantum discord Rulli and Sarandy (2011)), thereby enabling the same reduction in numerical cost. This suggests the potential applicability of our method to various measures of multipartite quantum correlation.
direct calculation efficient calculation
@@
infinite MPS
(trans. inv., normal)
@@
Table 1: Numerical costs for calculating the multipartite work deficit in $N$-site and infinite MPS systems up to constant multiples. Direct calculation scales exponentially with the system size $N$, making it infeasible for infinite systems. On the other hand, by employing the efficient calculation method introduced in this work, the costs are reduced to scale linearly with respect to $N$. For infinite systems described by translationally invariant normal MPS Cirac et al. (2017, 2021), this method achieves a constant cost, dependent only on the desired error bound $\varepsilon$, which scales as $e^{\mathcal{O}\left(\frac{1}{\varepsilon}\right)}$.
By utilizing our efficient method, we demonstrate the multipartite work deficit in several examples of quantum many-body systems. In the AKLT state Affleck et al. (1987, 1988) and the cluster state Briegel and Raussendorf (2001), representative examples of MPS, we analytically obtain the exact values for the multipartite work deficit. We also conduct numerical calculation in an MPS family Wolf et al. (2006), where we observe that the multipartite work deficit shows a sharp peak at the quantum phase transition point. This showcases that the multipartite work deficit serves as an efficiently computable probe of quantum phase transitions. We further show that the peak at the transition point becomes sharper under coarse-graining, which confirms that our probe indeed captures the singularity that is physically relevant (i.e. not from the artifact of optimization).
@@ II.1 Definition of multipartite work deficit
where $S(\rho)\equiv-\mathrm{Tr}[\rho\ln\rho]$ denotes the von Neumann entropy. This equation shows that the extractable work $W_{g}$ is characterized with the gap between the maximum entropy of the system $N\ln d$ and the initial entropy $S(\rho)$ of the total system. This can be understood as the consequence of the second law of thermodynamics, which states that a larger amount of work can be extracted from a state with high purity. The concrete protocols to achieve the work extraction in Eq. (2) have been studied in, e.g., Refs. Alicki et al. (2004); Jacobs (2009); Parrondo et al. (2015).
@@
This equation shows that $\Delta$ can be quantified with the optimization of the local measurements in $N$ parties, which can be exactly performed with reasonable numerical cost for small systems.
Figure 3: Schematic for extractable work through a measurement protocol $\Lambda$. Extractable work from $n$-th party $a_{n}$ is $\ln d-S(\tilde{\rho}_{n}^{\Lambda,y_{n}})$ when the measurement outcome $y_{n}$ is obtained, and the expectation value of the work extraction from $a_{n}$ becomes $\ln d-\sum_{y_{n}}p_{\Lambda}(y_{n})S(\tilde{\rho}_{n}^{\Lambda,y_{n}})$. On the other hand, the work required to erase the measurement outcomes $Y_{N}$ is $H_{\Lambda}(Y_{N})$. Therefore, the net work extraction can be quantified as the gap between these two contributions as shown in Eq. (3).
@@
We remark that multipartite work deficit is reduced to Eq. (6), which is equivalent to other measures of multipartite quantum correlations introduced in the previous works Piani et al. (2008); SaiToh et al. (2008); Modi et al. (2010); Yao et al. (2015); Braga et al. (2014) contexts totally different from thermodynamics. Here, we have formulated the multipartite work deficit based on thermodynamics. More importantly, we emphasize that the efficient calculation method discussed in Sec. III has not proposed in any of the previous works.
@@ II.3 Calculation in multipartite states composed of Bell pairs
The states composed of Bell pairs (9) and (10) are the cases that the equality of Eq. (11) is achieved. This can be shown from the fact that the right-hand side of the inequality (11) is reduced to the number of Bell pairs for these states, as shown in Figs. 5(d) and (e), and at the same time, these lower bounds can indeed be achieved with the aforementioned measurement protocols. We remark that in general multipartite states, the equalities of these inequalities are not satisfied due to the existence of genuinely multipartite quantum correlation. Therefore, in order to calculate work deficits in general many-body states, we need to go beyond these inequalities, which will be explained in the next section.
@@ III Efficient calculation in many-body systems
While the multipartite work deficit can be calculated exactly with reasonable numerical cost in few-body systems, the calculation in many-body systems is challenging due to the exponential growth of the numerical cost. The reason for such cost explosion is the necessity of optimizing all the local bases: we need to optimize $N$ sets of measurement bases in order to calculate the work deficit (5). This difficulty in optimizing local bases is not unique to the multipartite work deficit but is a common issue across various measures of multipartite quantum correlations Plenio and Vedral (2001); Wei et al. (2004); Barnum and Linden (2001); Wei and Goldbart (2003); Shimony (1995); Orús (2008); Shi et al. (2010); Orús and Wei (2010); Mousolou et al. (2013); Huang and Lin (2010); Orús et al. (2014); Shi et al. (2016); Orus and Wei (2010); Rulli and Sarandy (2011); Campbell et al. (2011, 2013); Power et al. (2015); Sun et al. (2015); Huang et al. (2018). In this section, we propose an efficient calculation method for the multipartite work deficit, which achieves significant cost reduction in the systems described with MPS (Table. 1).
@@ III.1 Overview of efficient calculation method
We first provide an overview of our efficient calculation method of the multipartite work deficit in many-body and infinite systems. Since the work deficit $\Delta$ increases proportionally with the number of party $N$ for typical $N$-partite states, we calculate the work deficit density, which is defined as
@@
This quantifies the leading term of the work deficit in many-body systems, and can remain finite even for the infinite systems. Because the work deficit density $\delta$ is a leading term of $\Delta$ defined with the optimization over the entire system (Eq. (5)), the numerical cost of directly calculating $\delta$ increases exponentially with the system size.
@@
Here, these bounds $U_{l}$ and $L_{l}$ are derived by dividing the entire system into $l$-site segments, and by performing optimization in these segments. The tightness of these bounds can be improved by increasing the locality parameter $l$, while the numerical costs are also increases with $l$. Especially for the systems described with MPS, we can derive the following inequality, which is also depicted in Fig. 6(a):
@@
This is one of the main results of this paper, along with Eq.(13). Since the computational costs for $U_{l}$ and $L_{l}$ are $Ne^{\mathcal{O}(l)}$ up to constant multiples as discussed in Sec. III.2, Eq. (14) implies that the evaluation of $\delta$ with an arbitrary accuracy $\varepsilon$ can be achieved at the cost that scales as $Ne^{\mathcal{O}(\frac{1}{\varepsilon})}$, as shown in Table.(1).
@@
where $A^{\sigma_{n}}$ is $D_{B}$ dimensional square matrices corresponding to the orthogonal basis |$\ket{\sigma_{n}}$ for $n$-th site, and $X$ is the matrix of the same size which represents the boundary condition (e.g., $X=\mathbbm{1}_{D_{B}}$ for periodic boundary condition). For such states, we can show that the numerical costs for calculating the lower and upper bounds are proportional to the system size $N$, which is much smaller than the exponential cost $e^{\mathcal{O}(N)}$ in the direct calculation. Furthermore, for the translationally invariant infinite system described with the normal MPS Cirac et al. (2021, 2017), these bounds can be calculated with the constant cost, by utilizing the translational invariance of the system. Normal MPS is a class of MPS with finite correlation length and represents the unique ground state of gapped one-dimensional local Hamiltonians Perez-Garcia et al. (2006). This cost reduction is particularly useful, compared to the direct calculation, which is not applicable to the infinite systems.

### v2

@@ Abstract
Quantum correlations are at the core of quantum mechanics and play a crucial role in various fields such as quantum information, condensed matter physics, high energy physics, and quantum thermodynamics. While bipartite quantum correlations have been extensively studied, multipartite quantum correlations in many-body systems remain elusive due to their complex structure. In particular, a primary challenge lies in the fact that the calculation of multipartite quantum correlation measure often requires exponential cost with respect to the system size. In this work, we tackle this problem by adopting a thermodynamic approach; we introduce a measure of multipartite quantum correlations based on the difference in extractable thermodynamic work by global operations and LOCC (local operations and classical communication). This can be regarded as a multipartite generalization of the work deficit, which has attracted attention as a thermodynamic measure of bipartite quantum correlation. A distinguishing feature of the thermodynamic approach to multipartite quantum correlation is that we can compare the degree of quantum correlations with clear operational meaning; for example, it is shown that W state has more quantum benefit in work extraction than GHZ state. Importantly, we develop an efficient calculation method of the multipartite work deficit. This efficient method works for a special class of quantum many-body systems described by MPS (matrix product states), where the numerical cost is shown to be proportional to the system size, significantly reducing the exponential cost required for the direct calculations. We demonstrate this efficient method in representative MPS such as the AKLT state and the cluster state, and analytically obtain the exact values of this measure. We further show that a quantum phase transition described by MPS is well captured by the multipartite work deficit, where a singular behavior appears and becomes sharper under coarse-graining, at the transition point. This shows that the multipartite work deficit does not only highlight the fundamental connection between multipartite quantum correlations and quantum thermodynamics, but also serves as an efficiently-computable probe of the structures of quantum many-body systems including their quantum phase transitions.
@@ I Introduction
In light of recent developments in quantum technologies Gross and Bloch (2017); Schäfer et al. (2020); Kjaergaard et al. (2020); Krantz et al. (2019); Lukin et al. (2020); Altman et al. (2021), quantum correlations are becoming increasingly important not only in quantum computation and quantum communication Gisin and Thew (2007); Nielsen and Chuang (2010); Briegel et al. (2009) but also in various fields of physics including condensed matter physics Amico et al. (2008); De Chiara and Sanpera (2018) and high energy physics Ryu and Takayanagi (2006a, b); Takayanagi (2012). For example, quantum correlations have been utilized as useful probes for understanding critical phenomena Vidal et al. (2003); Calabrese and Cardy (2004, 2009), topological order Kitaev and Preskill (2006); Levin and Wen (2006), and information scrambling Swingle et al. (2016). While the understanding on bipartite quantum correlations has been well-established through various measures with clear operational meanings Plenio and Virmani (2005); Horodecki et al. (2009); Modi et al. (2012); Bera et al. (2017), multipartite quantum correlations remain a very active research avenue Plenio and Vedral (2001); Wei et al. (2004); Barnum and Linden (2001); Wei and Goldbart (2003); Shimony (1995); Orús (2008); Shi et al. (2010); Orús and Wei (2010); Huang and Lin (2010); Azimi Mousolou et al. (2013); Orús et al. (2014); Shi et al. (2016); Orus and Wei (2010); Rulli and Sarandy (2011); Campbell et al. (2011, 2013); Power et al. (2015); Sun et al. (2015); Huang et al. (2018); Braga et al. (2012); Xu (2013); Liu et al. (2014); Coffman et al. (2000); Radhakrishnan et al. (2020); Hu et al. (2018); Hyllus et al. (2012); Tóth (2012); Fröwis and Dür (2012); Hauke et al. (2016); Mermin (1990); Ardehali (1992); Pezze et al. (2017), mainly because of their rich and complicated structure Gühne and Tóth (2009); Bengtsson and Zyczkowski (2016). For example, even in tripartite quantum states, there are different types of quantum correlations represented by the W state and the GHZ state.
@@
Furthermore, we develop an efficient calculation method of the multipartite quantum correlations for a special class of quantum many-body systems described by MPS (matrix product states) Perez-Garcia et al. (2006); Verstraete et al. (2008); Cirac et al. (2017, 2021); Orús (2014). While the cost of naive direct calculation of the multipartite work deficit grows exponentially with the system size as is the case for many other measures of multipartite quantum correlations Plenio and Vedral (2001); Wei et al. (2004); Barnum and Linden (2001); Wei and Goldbart (2003); Shimony (1995); Orús (2008); Shi et al. (2010); Orús and Wei (2010); Azimi Mousolou et al. (2013); Huang and Lin (2010); Orús et al. (2014); Shi et al. (2016); Orus and Wei (2010); Rulli and Sarandy (2011); Campbell et al. (2011, 2013); Power et al. (2015); Sun et al. (2015); Huang et al. (2018), the cost of our efficient method is proportional to the system size enabling the significant cost reduction (Table 1). It is noteworthy that this efficient method is not restricted to the work deficit but also is applicable to another measure of multipartite quantum correlation (i.e., global quantum discord Rulli and Sarandy (2011)), thereby enabling the same reduction in numerical cost. This suggests the potential applicability of our method to various measures of multipartite quantum correlation.
Direct calculation
Efficient calculation
with error bound $\varepsilon$
@@
Normal MPS in
thermodynamic limit
@@
Table 1: Numerical costs for calculating the multipartite work deficit in $N$-site MPS and MPS in the thermodynamic limit, up to constant multiples. Direct calculation scales exponentially with the system size $N$, making it infeasible in the thermodynamic limit. On the other hand, by employing the efficient calculation method introduced in this work, the costs are reduced to scale linearly with respect to $N$. Furthermore, this method is applicable even in the thermodynamic limit, for translationally invariant normal MPS Cirac et al. (2017, 2021), achieving a constant cost dependent only on the desired error bound $\varepsilon$, which scales as $e^{\mathcal{O}\left(\frac{1}{\varepsilon}\right)}$.
By utilizing our efficient method, we demonstrate the multipartite work deficit in several examples of quantum many-body systems. In the AKLT state Affleck et al. (1987, 1988) and the cluster state Briegel and Raussendorf (2001), representative examples of MPS, we analytically obtain the exact values for the multipartite work deficit. We also conduct numerical calculation in an MPS family Wolf et al. (2006), where we observe that the multipartite work deficit shows a sharp kink at the quantum phase transition point. This showcases that the multipartite work deficit serves as an efficiently computable probe of quantum phase transitions. We further show that the kink at the transition point becomes sharper under coarse-graining, which confirms that our probe indeed captures the singularity that is physically relevant (i.e. not from the artifact of optimization).
@@ II.1 Definition of multipartite work deficit
where $S(\rho)\equiv-\mathrm{Tr}[\rho\ln\rho]$ denotes the von Neumann entropy. This equation shows that the extractable work $W_{g}$ is characterized with the gap between the maximum entropy of the system $N\ln d$ and the entropy $S(\rho)$ of the total system. This can be understood as the consequence of the second law of thermodynamics, which states that a larger amount of work can be extracted from a state with high purity. The concrete protocols to achieve the work extraction in Eq. (2) have been studied in, e.g., Refs. Alicki et al. (2004); Jacobs (2009); Parrondo et al. (2015).
@@
This equation shows that $\Delta$ is quantified with the optimization of the local measurements on $N$ parties $\Lambda\in\mathcal{M}_{N}$, which can be exactly performed with reasonable numerical cost for small systems.
Figure 3: Schematic for extractable work through a measurement protocol $\Lambda$. Extractable work from $n$-th party $a_{n}$ is $\ln d-S(\tilde{\rho}_{n}^{\Lambda,y_{n}})$ when the measurement outcome $y_{n}$ is obtained, and the averaged work extraction becomes $\ln d-\sum_{y_{n}}p_{\Lambda}(y_{n})S(\tilde{\rho}_{n}^{\Lambda,y_{n}})$. On the other hand, the work required to erase the measurement outcomes $Y_{N}$ is $H_{\Lambda}(Y_{N})$. Therefore, the net work extraction can be quantified as the gap between these two contributions as shown in Eq. (3).
@@
We remark that multipartite work deficit is reduced to Eq. (6), which is equivalent to other measures of multipartite quantum correlations introduced in the previous works Piani et al. (2008); SaiToh et al. (2008); Modi et al. (2010); Yao et al. (2015); Braga et al. (2014), in the contexts totally different from thermodynamics. Here, we have formulated the multipartite work deficit based on thermodynamics. More importantly, we emphasize that the efficient calculation method discussed in Sec. III has not proposed in any of the previous works.
@@ II.3 Calculation in multipartite states composed of Bell pairs
The states composed of Bell pairs (9) and (10) are the cases that the equality of Eq. (11) is achieved. This can be shown from the fact that the right-hand side of the inequality (11) is reduced to the number of Bell pairs for these states, as shown in Figs. 5(d) and (e), and at the same time, these lower bounds can indeed be achieved with the aforementioned measurement protocols. We remark that in general multipartite states, the equality of Eq. (11) is not satisfied due to the existence of genuinely multipartite quantum correlation. Therefore, in order to calculate work deficits in general many-body states, we need to go beyond these inequalities, which will be explained in the next section.
@@ III Efficient calculation in many-body systems
While the multipartite work deficit can be calculated exactly with a reasonable numerical cost in few-body systems, the calculations in many-body systems are challenging due to the exponential growth of the numerical cost. The reason for such a cost explosion is the necessity of optimizing all the local bases: we need to optimize $N$ sets of measurement bases in order to calculate the work deficit (5), which requires an exponential cost in the brute-force optimization. Such a difficulty in optimizing local bases is not unique to the multipartite work deficit but is a common issue across various measures of multipartite quantum correlations Plenio and Vedral (2001); Wei et al. (2004); Barnum and Linden (2001); Wei and Goldbart (2003); Shimony (1995); Orús (2008); Shi et al. (2010); Orús and Wei (2010); Azimi Mousolou et al. (2013); Huang and Lin (2010); Orús et al. (2014); Shi et al. (2016); Orus and Wei (2010); Rulli and Sarandy (2011); Campbell et al. (2011, 2013); Power et al. (2015); Sun et al. (2015); Huang et al. (2018). While there have been some attempts to overcome this problem  [^1] In some previous works Orús (2008); Shi et al. (2010); Orús and Wei (2010); Azimi Mousolou et al. (2013); Huang and Lin (2010); Orus and Wei (2010); Sun et al. (2015); Huang et al. (2018), the numerical costs are reduced by posing assumptions on the optimal local bases. For example, in the translationally invariant systems, the optimal bases are assumed to be also translationally invariant. While such an assumption is powerful in reducing the numerical cost and also seems to be natural, there is no solid theoretical guarantee on it, to the best of our knowledge., systematic methods to calculate these measures with polynomial costs are yet to be established. In this section, we propose an efficient calculation method for the multipartite work deficit, which achieves significant cost reduction in the systems described with MPS (Table 1). We note that this efficient calculation method is applicable to another measure of multipartite quantum correlation called global quantum discord Rulli and Sarandy (2011), as detailed in Supplemental Material SM .
@@ III.1 Overview of efficient calculation method
We first provide an overview of our efficient calculation method of the multipartite work deficit in many-body systems. Since the work deficit $\Delta$ increases proportionally with the number of party $N$ for typical $N$-partite states, we focus on the work deficit density, which is defined as
@@
This quantifies the leading term of the work deficit in many-body systems, and does not diverge even in the thermodynamic limit. Because the work deficit density $\delta$ is a leading term of $\Delta$ defined with the optimization over the entire system (Eq. (5)), the numerical cost of directly calculating $\delta$ increases exponentially with the system size.
@@
Here, these bounds $U_{l}$ and $L_{l}$ are derived by dividing the entire system into $l$-site segments, and by performing optimization in these segments. The tightness of these bounds can be improved by increasing the locality parameter $l$, while the numerical costs also increase with $l$. Especially for the systems described with MPS, we can derive the following inequality, which is also depicted in Fig. 6(a):
@@
This is one of the main results of this paper, along with Eq.(13). Since the computational costs for $U_{l}$ and $L_{l}$ are $Ne^{\mathcal{O}(l)}$ up to constant multiples as discussed later in Sec. III.2, Eq. (14) implies that the evaluation of $\delta$ with an arbitrary accuracy $\varepsilon$ can be achieved at the cost that scales as $Ne^{\mathcal{O}(\frac{1}{\varepsilon})}$, as shown in Table 1.
@@
where $A^{\sigma_{n}}$ is $D_{B}$-dimensional square matrix corresponding to the orthogonal basis |$\ket{\sigma_{n}}$ for $n$-th site, and $X$ is the matrix of the same size which represents the boundary condition (e.g., $X=\mathbbm{1}_{D_{B}}$ for periodic boundary condition and $X=\bm{v}_{R}\bm{v}_{L}^{T}$ for open boundary condition, with certain $D_{B}$-dimensional vectors $\bm{v}_{R}$ and $\bm{v}_{L}$). For such states, the numerical costs for calculating the lower and upper bounds are proportional to the system size $N$, which is much smaller than the exponential cost $e^{\mathcal{O}(N)}$ in the direct calculation. Furthermore, these bounds are calculable even in the thermodynamic limit for translationally invariant normal MPS Cirac et al. (2021, 2017), where the direct calculation of $\delta$ is infeasible. Normal MPS is a class of MPS with finite correlation length and represents the unique ground state of gapped one-dimensional local Hamiltonians Perez-Garcia et al. (2006).